peform the indicated operation
3/8+(-4/9)

Answer by nyc_function(2741)

3/8 +(-4/9) =


3/8 - 4/8 =


LCD = 72


Applying the rules of subtraction of fractions, we get:


(27 - 4(8))/72 =


(27 - 32)/72 = -5/72


Done!